OpenStudy (anonymous):

Math model problem please show work will give you a medal (6×5+2)-(4×5-3)= A.15 B.21 C.9

OpenStudy (anonymous):

A. 15 @flores210am

OpenStudy (anonymous):

6x5=30 +2=32 4x5=20 -3=17 17-32=

OpenStudy (anonymous):

15